OpenStudy (anonymous):

(x-2)^2=100 Solve the quadratic function with square roots

OpenStudy (tkhunny):

It's already nicely set up for you. Find the two square roots and you're gold.

OpenStudy (anonymous):

would the answer be 12?

OpenStudy (tkhunny):

That's one. Where did the other one go? It helps to write out your process, not just your answer.

OpenStudy (anonymous):

theres another one?

OpenStudy (tkhunny):

Two. Hint: (2)^2 = 4 (-2)^2 = 4

OpenStudy (xapproachesinfinity):

if \[x^2=a \Longrightarrow |x|=\sqrt{a}\]

OpenStudy (danjs):

\[\large \sqrt{u^2}=\left| u \right|\]

OpenStudy (danjs):

even power

OpenStudy (xapproachesinfinity):

the reasoning is given by @tkhunny when you square positive and number like (-2)^2 (2)^2 the result is the same 4

OpenStudy (xapproachesinfinity):

that is why we are using absolute value

OpenStudy (anonymous):

If I squared (x-2) all together and 100, I would get x-2=10 correct?

OpenStudy (tkhunny):

Well, after all that, did you find the second one? Here's how you find the first one: (x-2)^2 = 100 Square Root x-2 = 10 ==> x = 12 What you did not do was this: Square Root x-2 = -10 ==> What?

OpenStudy (danjs):

yes, if you square root something that is squared like root of (x-2)^2, that something can be positive or negative, \[\sqrt{(x-2)^2} = \left| x-2 \right| = 10\]

OpenStudy (danjs):

solving absolute value probs, that x-2 can be +10 or -10 2 values for x as @tkhunny showed above